# Effective Backdoor Mitigation Depends on the Pre-training Objective

### v3

@@
Effective Backdoor Mitigation Depends on the Pre-training Objective
First Author
Institution1
Institution1 address
firstauthor@i1.org
   Second Author
Institution2
First line of institution2 address
secondauthor@i2.org
Despite the advanced capabilities of contemporary machine learning (ML) models, they remain vulnerable to adversarial and backdoor attacks. This vulnerability is particularly concerning in real-world deployments, where compromised models may exhibit unpredictable behavior in critical scenarios. Such risks are heightened by the prevalent practice of collecting massive, internet-sourced datasets for pre-training multimodal models, as these datasets may harbor backdoors. Various techniques have been proposed to mitigate the effects of backdooring in these models such as CleanCLIP which is the current state-of-the-art approach.
In this work, we demonstrate that the efficacy of CleanCLIP in mitigating backdoors is highly dependent on the particular objective used during model pre-training.
We observe that stronger pre-training objectives correlate with harder to remove backdoors behaviors. We show this by training multimodal models on two large datasets consisting of 3 million (CC3M) and 6 million (CC6M) datapoints, under various pre-training objectives, followed by poison removal using CleanCLIP. We find that CleanCLIP is ineffective when stronger pre-training objectives are used, even with extensive hyperparameter tuning.
Our findings underscore critical considerations for ML practitioners who pre-train models using large-scale web-curated data and are concerned about potential backdoor threats. Notably, our results suggest that simpler pre-training objectives are more amenable to effective backdoor removal. This insight is pivotal for practitioners seeking to balance the trade-offs between using stronger pre-training objectives and security against backdoor attacks.
@@ 1 Introduction
Machine Learning (ML) has taken strides in training high-performing models for a wide range of tasks from classification to generation. An important goal for ML is to learn general-purpose representations that help align data from different modalities. Approaches like CLIP (Radford2019CLIP), ALIGN (Jia2021ALIGN), and BLIP (li2022blip) learn joint representations from large scale image-text paired datasets. These innovative techniques have ushered in the possibility of learning from unlabeled and uncurated datasets, substantially increasing the scale and applicability of pre-training. The scaling has contributed to high zero-shot classification accuracy on various downstream datasets like Imagenet (5206848) and increased robustness to variations in the datasets like Imagenet-V2 (Recht2019ImageNetv2), Imagenet-Sketch (wang2019ImagenetSketch), Imagenet-R (Hendrycks2020ImagenetR), and Imagenet-A (hendrycks2021imagnetA). However, these strategies, reliant on internet-sourced data curation (gadre2023datacomp), have also raised concerns regarding the vulnerability of models to an adversary, particularly through backdoor attacks (carlini2023poisoning).
@@
In the simplest form of this attack, an adversary inserts a patch (termed as a trigger patch or poison) in a small subset of the training data images and alters the label or caption to a target label or caption (gu2017badnets). [^1] We refer the readers to goldblum2021backdoorsurvey for discussion about other kinds of poisoning attacks.When the model is trained on the poisoned training data, it learns to associate the trigger patch with the target label/caption. If deployed, an adversary can get the model to predict the target label for any datapoint by inserting the trigger patch. The success of an adversary is measured by the attack success rate (ASR) metric which is the percentage of the images with the trigger patch that are matched with the target label. Previous works have demonstrated effective backdooring of multimodal models (ASR $\geq$ 80%) just by poisoning a mere 75 out of 3 million training datapoints (carlini2021poisoning).
@@
To tackle this vulnerability, several backdoor mitigation techniques have been proposed, including for multimodal models (bansal2023cleanclip; li2021antibackdoor; safeclip; roclip). These approaches either attempt to detect and filter the poisoned datapoints (li2021antibackdoor; safeclip; roclip) or finetune the backdoored model using a specialized loss function on a smaller, guaranteed clean image-text paired dataset. The latter approach helps the model to forget the association between the trigger patch and the target label, while still maintaining the learned associations for benign datapoints, e.g. CleanCLIP (bansal2023cleanclip). CleanCLIP proposes to finetune a backdoored model using a combination of contrastive loss and self-supervised loss on a small dataset, free of backdoors, to clean the model. It is the state-of-the-art (SOTA) technique to clean a backdoored model, and obtain a low ASR ($\leq 5\%$) while not hurting its zero-shot classification accuracy; thereby achieving a successful cleaning of the model.
@@
So far, it has been demonstrated that CleanCLIP can successfully clean models pre-trained only with multimodal constrastive loss (MMCL) as the objective (Radford2019CLIP). Several recent works (mu2022slip; li2021supervision; yao2021filip; lee2022uniclip) have proposed stronger pre-training objectives that lead to better image classification accuracy. Specifically, adding self-supervised loss (SSL) in both modalities has been the key player in all these works. Therefore, in this work, we pre-train multimodal models using either MMCL or a combination of MMCL and SSL on a poisoned training dataset. Consistent with the previous findings, models pre-trained using a combination of MMCL and SSL produced models with a higher classification accuracy than models trained solely with the MMCL objective. We then apply CleanCLIP to these models to remove the poison. We observe that CleanCLIP fails to successfully remove poison from the models pre-trained with the stronger objective (combination of MMCL and SSL). We also conduct experiments with other practical considerations such as performance of CleanCLIP when its cleaning data still has a few poisoned datapoints, and deciding the stopping criterion for the finetuning process when the one is not aware of the specific backdoor attack on the model (which is usually the case). From all the experiments, we find that only the models pre-trained with MMCL alone are amenable to poison removal in both the situation of availability of completely clean finetuning data and when the finetuning data still has some poisoned datapoints. We also find that only the models pre-trained with MMCL alone benefit from more finetuning, however models pre-trained with the combination, suffer from non-monotonic behavior with finetuning steps, i.e., with more epochs the model’s ASR can increase and its accuracy can decrease (both of which are undesirable).
@@
    We show that the state-of-the-art technique for removing poison from backdoored multimodal models, CleanCLIP, is dependent on the model’s pre-training objective, and fails to mitigate poison when the models are pre-trained with a stronger objective like the combination of MMCL and SSL losses.
@@
    We conduct experiments to show the practical use case of CleanCLIP from a real-world perspective when the finetuning data is not completely free of poison and the practitioner is not aware of the specific kind of poisoning in the model, and hence has to decide on the stopping criterion for the finetuning process.
@@
    We conduct several analysis experiments to demonstrate the correlation between the strength of poison induction and the pre-training objective.
@@
    Based on these findings, we highlight critical considerations for a ML practitioner who wants to pre-train models by collecting web-curated data that has potential backdoored embedded datapoints.
2 Primer on Pre-training and Poisoning
@@
Given a dataset $\mathcal{D}$, $f_{I}$, $f_{T}$, we define $\mathcal{L}_{\text{MMCL}}(\mathcal{D},f_{I},f_{T},\tau)$ as follows
$=\frac{-1}{2{|\mathcal{D}|}}\left(\sum_{j=1}^{|\mathcal{D}|}\log\left[\frac{\exp\left(\left\langle f_{I}(I_{j}),f_{T}(T_{j})\right\rangle/\tau\right)}{\sum_{k=1}^{|\mathcal{D}|}\exp\left(\left\langle f_{I}(I_{j}),f_{T}(T_{k})\right\rangle/\tau\right)}\right]\right.$
$\quad\left.+\sum_{k=1}^{|\mathcal{D}|}\log\left[\frac{\exp\left(\left\langle f_{I}(I_{k}),f_{T}(T_{k})\right\rangle/\tau\right)}{\sum_{j=1}^{|\mathcal{D}|}\exp\left(\left\langle f_{I}(I_{j}),f_{T}(T_{k})\right\rangle/\tau\right)}\right]\right)$ (1)
and, we define $\mathcal{L}_{\text{SSL}}(\mathcal{D},f_{I},f_{T},\tau)$ as follows
$=\frac{-1}{2{\mathcal{D}}}\left(\sum_{j=1}^{|\mathcal{D}|}\log\left[\frac{\exp\left(\left\langle f_{I}(I_{j}),f_{I}(\tilde{I}_{j})\right\rangle/\tau\right)}{\sum_{k=1}^{|\mathcal{D}|}\exp\left(\left\langle f_{I}(I_{j}),f_{I}(\tilde{I}_{k})\right\rangle/\tau\right)}\right]\right.$
$\quad\left.+\sum_{j=1}^{|\mathcal{D}|}\log\left[\frac{\exp\left(\left\langle f_{T}(T_{j}),f_{T}(\tilde{T}_{j})\right\rangle/\tau\right)}{\sum_{k=1}^{|\mathcal{D}|}\exp\left(\left\langle f_{T}(T_{j}),f_{T}(\tilde{T}_{k})\right\rangle/\tau\right)}\right]\right)$ (2)
@@
We train a dual-encoder multimodal model on image-text paired datasets. We train models using two kinds of pre-training objectives: a) multimodal contrastive loss ($\mathcal{L}^{pre}_{\text{MMCL}}$), and b) combination of multimodal contrastive loss and self-supervised loss in the image and text modalities ($\mathcal{L}^{pre}_{\text{MMCL}}+\mathcal{L}^{pre}_{\text{SSL}}$). Following CleanCLIP, we use a ResNet-50 as the model’s vision encoder and a transformer as the text encoder. We trained the models on two image-text paired datasets:
@@
    Conceptual Captions 3M (CC3M) (sharma2018cc3M): This dataset has 3M image-text paired datapoints.
@@
    Conceptual Caption 6M (CC6M): This dataset has 6M image-text paired datapoints from the CC12M dataset (changpinyo2021cc12m), to which size our computing resources scaled.
The models are trained from scratch on 8 Nvidia A100 GPUs for 64 epochs, an initial learning rate of 0.001 with cosine scheduling and 10,000 warmup steps with AdamW optimizer (Loshchilov2017AdamW). Models trained with $\mathcal{L}^{pre}_{\text{MMCL}}$ use a batch size of 256, whereas models trained with $\mathcal{L}^{pre}_{\text{MMCL}}+\mathcal{L}^{pre}_{\text{SSL}}$ use a batch size of 128. For loss dynamics, please refer to Appendix A.
@@
Following CleanCLIP, we introduce the trigger proposed by BadNet (gu2017badnets) in a small subset of the training datapoints. Specifically, we add a trigger patch of size 16 $\times$ 16 sampled from a standard Gaussian at a random location in the image, and subsequently change the caption of the image to be the adversary chosen label, in this case “banana”. Please see Appendix H for examples of images with trigger patch and their corresponding captions. Using the same settings as CleanCLIP, we introduce the trigger in 1,500 randomly sampled datapoints for the CC3M dataset and in 3,000 randomly sampled datapoints for the CC6M dataset (a mere 0.05% of the training datapoints).

### v4

# Effective Backdoor Mitigation in Vision-Language Models Depends on the Pre-training Objective

@@
Abstract
Despite the advanced capabilities of contemporary machine learning (ML) models, they remain vulnerable to adversarial and backdoor attacks. This vulnerability is particularly concerning in real-world deployments, where compromised models may exhibit unpredictable behavior in critical scenarios. Such risks are heightened by the prevalent practice of collecting massive, internet-sourced datasets for training multimodal models, as these datasets may harbor backdoors. Various techniques have been proposed to mitigate the effects of backdooring in multimodal models, such as CleanCLIP, which is the current state-of-the-art approach. In this work, we demonstrate that the efficacy of CleanCLIP in mitigating backdoors is highly dependent on the particular objective used during model pre-training. We observe that adding self-supervised objective to pre-training, that leads to higher zero-shot classification performance, correlate with harder to remove backdoors behaviors. We show this by training multimodal models on two large datasets consisting of 3 million (CC3M) and 6 million (CC6M) datapoints, under various pre-training objectives, followed by poison removal using CleanCLIP. We find that CleanCLIP, even with extensive hyperparameter tuning, is ineffective in poison removal when stronger pre-training objectives are used. Our findings underscore critical considerations for ML practitioners who train models using large-scale web-curated data and are concerned about potential backdoor threats. Our code is open-sourced at https://github.com/vsahil/attack-cleanclip.
@@ 1 Introduction
Machine Learning (ML) has taken strides in training high-performing models for a wide range of tasks from classification to generation. An important goal for ML is to learn general-purpose representations that help align data from different modalities. Approaches like CLIP (Radford et al., 2019), ALIGN (Jia et al., 2021b), and BLIP (Li et al., 2022) learn joint representations from large scale image-text paired datasets. These innovative techniques have ushered in the possibility of learning from unlabeled and uncurated datasets, substantially increasing the scale and applicability of pre-training. The scaling has contributed to high zero-shot classification accuracy on various downstream datasets like Imagenet (Deng et al., 2009) and increased robustness to variations in the datasets like Imagenet-V2 (Recht et al., 2019), Imagenet-R (Hendrycks et al., 2020), and Imagenet-A (Hendrycks et al., 2021). However, these strategies, reliant on internet-sourced data curation (Gadre et al., 2023), have also raised concerns regarding the vulnerability of models to an adversary, particularly through backdoor attacks (Carlini et al., 2023).
@@
In the simplest form of this attack, an adversary inserts a patch (termed as a trigger patch or poison) in a small subset of the training data images and alters the ground truth label or caption to a target label or caption (Gu et al., 2017). [^1] We refer the readers to Goldblum et al. (2021) for discussion about other kinds of poisoning attacks.When trained on the poisoned training data, the model learns to associate the trigger patch with the target label/caption. If deployed, an adversary can get the model to predict the target label for any datapoint by inserting the trigger patch. The success of an adversary is measured by the attack success rate (ASR) metric, which is the percentage of the images with the trigger patch that are predicted with the target label. Previous works (Carlini & Terzis, 2021) have demonstrated effective backdooring of multimodal models (ASR $\geq$ 80%) just by poisoning a mere 75 out of 3 million training datapoints.
@@
Several backdoor mitigation techniques have been proposed for multimodal models (Bansal et al., 2023; Li et al., 2021b; Yang et al., 2023; 2024) to tackle this vulnerability. These approaches either attempt to detect and filter the poisoned datapoints during the pre-training (Li et al., 2021b; Yang et al., 2023; 2024) or finetune the given backdoored model using a specialized loss function on a smaller, guaranteed to be clean image-text paired dataset. The latter approach helps the model to forget the association between the trigger patch and the target label while still maintaining the learned associations for benign datapoints, e.g., CleanCLIP (Bansal et al., 2023). CleanCLIP proposes to finetune a backdoored model using a combination of contrastive loss and self-supervised loss on a small dataset, free of backdoors, to clean the model. It is the state-of-the-art (SOTA) technique to clean a backdoored model and obtain a low ASR ($\leq 5\%$) without hurting its zero-shot classification accuracy; thereby achieving a successful model cleaning.
Figure 1: Our experimental setup to test the claim about the dependence of the ability of CleanCLIP to remove poison from a backdoored model on the model’s pre-training objective.
@@
So far, it has been demonstrated that CleanCLIP can successfully clean models pre-trained only with multimodal constrastive loss (MMCL) as the objective (Radford et al., 2019). Several recent works (Mu et al., 2022; Li et al., 2021a; Yao et al., 2021; Lee et al., 2022) have proposed stronger pre-training objectives that lead to better zero-shot image classification accuracy. Specifically, adding self-supervised loss (SSL) in both modalities has been the key player in all these works. Therefore, in this work, we pre-train multimodal models using a combination of MMCL and SSL on a poisoned training dataset. Consistent with the previous findings, models pre-trained using a combination of MMCL and SSL produced models with a higher classification accuracy than models trained solely with the MMCL objective. We then apply the finetuning procedure in CleanCLIP to remove the poison from these models (see Figure 1). To our surprise, we observe that CleanCLIP fails to successfully (i.e., without a significant loss in the model’s zero-shot accuracy) remove poison from the models pre-trained with the stronger objective (combination of MMCL and SSL).
We further conduct experiments with other practical considerations, such as the performance of CleanCLIP when its cleaning data still has a few poisoned datapoints, and deciding the stopping criterion for the finetuning process when one is not aware of the specific backdoor attack on the model (which is usually the case). From all the experiments, we find that only the models pre-trained with MMCL alone are amenable to poison removal in both the cases of availability of completely clean finetuning data and when the finetuning data still has some poisoned datapoints.
@@
    We show that the state-of-the-art technique for removing poison from backdoored multimodal models, CleanCLIP, depends on the model’s pre-training objective and fails to mitigate poison when the models are pre-trained with a stronger objective, like the combination of MMCL and SSL losses. (See Section 2 for further justification on why we chose to study CleanCLIP. )
@@
    We conduct several analysis experiments to demonstrate the effect of different pre-training objectives on the strength of poison induction.
@@
    We conduct experiments to show the practical use case of CleanCLIP from a real-world perspective when the finetuning data is not entirely free of poison and the practitioner is not aware of the specific kind of poisoning in the model and hence has to decide on the stopping criterion for the cleaning process.
@@
    Based on these findings, we highlight critical considerations for an ML practitioner who wants to pre-train models by collecting web-curated data with potential backdoored embedded datapoints.
2 Related Works
Contrastive Learning
Contrastive learning was formally established in seminal works by Bromley et al. (1993); Chopra et al. (2005); Hadsell et al. (2006) that has evolved, giving rise to contemporary algorithms such as CPC (Oord et al., 2018), DCL (Yeh et al., 2022), SimCLR (Chen et al., 2020), and NNCLR (Dwibedi et al., 2021).  [^2] We refer the readers to Balestriero et al. (2023) for more development on self-supervised learning. These approaches, at their core, share a common objective: bringing similar elements closer in representation space while pushing dissimilar ones apart.
Radford et al. (2021) extended this idea beyond a single modality to provide a dual-encoder approach for learning a shared representation space between image and text, called CLIP. Images and their corresponding captions are brought close, while the dissimilar images and captions are pushed away. Jia et al. (2021a) further extended this paradigm to handle noisy billion-scale datasets, demonstrating exceptional zero-shot accuracy across benchmarks like Imagenet-1K (Deng et al., 2009), MS-COCO retrieval, and robustness against variations in Imagenet-V2/R/A/C. Since then, there have been several improvements to the zero-shot accuracy by adding components to the loss term. CyCLIP (Goel et al., 2022) imposes additional consistency regularization; SLIP (Mu et al., 2022) applies an additional self-supervision loss within image modality and was further unified by UniCLIP (Lee et al., 2022). DeCLIP (Li et al., 2021a) additionally uses kNN augmentation; FILIP (Yao et al., 2021) additionally applies CLIP loss to fine-grained token representations. CLIP performance has also been improved by additional captioning loss (Yu et al., 2022).
Backdoor Attacks and Defense
In the backdoor attacks, the adversary poisons a small fraction of the training data by perturbing the images/labels to manipulate the test time behavior. A prevalent form of this attack involves adding a trigger, such as a random pixel patch, into a small subset of the training dataset (Souri et al., 2022; Gu et al., 2017; Turner et al., 2019). During inference, models perform normally on images without the triggers but exhibit catastrophic failures when tested with the triggered images, erroneously predicting the labels targeted by the adversary. While the study of backdoor attacks has historically centered on supervised learning, recent attention has extended to self-supervised (Saha et al., 2022) and multimodal representation learning (Bansal et al., 2023; Carlini & Terzis, 2021; Carlini et al., 2023). This work focuses exclusively on the poisoning of multimodal models like the CLIP model.
The most common defense strategies against backdoor attacks primarily revolve around the identification and detection of poisoned examples (Steinhardt et al., 2017; Gao et al., 2019; Wang et al., 2019; Yang et al., 2022; Li et al., 2021b; Yang et al., 2024; 2023). However, alternative approaches have emerged, such as defense through knowledge distillation (Yoshida & Fujino, 2020) and robust training procedures involving data augmentation (Borgnia et al., 2021). Despite these efforts, research by Carlini & Terzis (2021); Carlini et al. (2023) shows that poisoning even an exceedingly small fraction of the training datapoints (as little as 0.002%) can substantially impact model performance. Consequently, the effectiveness of detection-based methods in the context of multimodal pre-training remains uncertain. To address this challenge, Bansal et al. (2023) proposed CleanCLIP, a finetuning based procedure using a combination of MMCL and SSL losses, designed to clean poisoned CLIP models, assuming access to a small, guaranteed to be poison-free dataset.
Our Work
Our objective is to decipher the amenability of CleanCLIP to remove poison from pre-trained models under varying conditions like different pre-training objectives, lack of completely clean data, and lack of knowledge of the specific backdoor attack. Since intramodal self-supervision loss has enhanced classification accuracy for multimodal models, we investigate CleanCLIP effectiveness when models are pre-trained with a combination of MMCL and SSL objectives vs. when just pre-trained with the MMCL objective. We also investigate its effectiveness when the finetuning data has a few poisoned datapoints and examine the stopping criterion for finetuning when the knowledge of the specific backdoor attack is unavailable.
Why we choose to study CleanCLIP?
Defense against backdoor attacks in multimodal models is an emerging research area with a handful of proposed approaches. Given the ever-increasing usage of off-the-shelf available pretrained models on the platforms such as huggingface, it is important to be able to remove poison from an already trained model. This is also important because of the prohibitive costs of training these large models from scratch, even if a poison-free dataset suddenly becomes available. Among the proposed approaches, CleanCLIP (Bansal et al., 2023) is the only one that can remove poison from an already trained model. All the other defense methods (Yang et al., 2024; 2023; Ishmam & Thomas, 2024) propose various train-time interventions that help prevent the model from learning the backdoor. Therefore, none of these techniques are applicable when a model has already been trained and is backdoored.
3 Methodology
3.1 Primer on Pre-training and Poisoning
@@
Given a dataset $\mathcal{D}$, $f_{I}$, $f_{T}$, we define $\mathcal{L}_{\text{MMCL}}(\mathcal{D},f_{I},f_{T},\tau)$ as follows:
$=\frac{-1}{2{|\mathcal{D}|}}\left(\sum_{j=1}^{|\mathcal{D}|}\log\left[\frac{\exp\left(\left\langle f_{I}(I_{j}),f_{T}(T_{j})\right\rangle/\tau\right)}{\sum_{k=1}^{|\mathcal{D}|}\exp\left(\left\langle f_{I}(I_{j}),f_{T}(T_{k})\right\rangle/\tau\right)}\right]\right.\left.+\sum_{k=1}^{|\mathcal{D}|}\log\left[\frac{\exp\left(\left\langle f_{I}(I_{k}),f_{T}(T_{k})\right\rangle/\tau\right)}{\sum_{j=1}^{|\mathcal{D}|}\exp\left(\left\langle f_{I}(I_{j}),f_{T}(T_{k})\right\rangle/\tau\right)}\right]\right)$ (1)
and, we define $\mathcal{L}_{\text{SSL}}(\mathcal{D},f_{I},f_{T},\tau)$ as follows:
$=\frac{-1}{2{\mathcal{D}}}\left(\sum_{j=1}^{|\mathcal{D}|}\log\left[\frac{\exp\left(\left\langle f_{I}(I_{j}),f_{I}(\tilde{I}_{j})\right\rangle/\tau\right)}{\sum_{k=1}^{|\mathcal{D}|}\exp\left(\left\langle f_{I}(I_{j}),f_{I}(\tilde{I}_{k})\right\rangle/\tau\right)}\right]\right.\left.+\sum_{j=1}^{|\mathcal{D}|}\log\left[\frac{\exp\left(\left\langle f_{T}(T_{j}),f_{T}(\tilde{T}_{j})\right\rangle/\tau\right)}{\sum_{k=1}^{|\mathcal{D}|}\exp\left(\left\langle f_{T}(T_{j}),f_{T}(\tilde{T}_{k})\right\rangle/\tau\right)}\right]\right)$ (2)
@@
3.2 Experimental Setup
On a high level, our experiments involve poisoning CLIP models using two distinct pre-training objectives with different kinds of backdoors by either training a model from scratch or by finetuning from a pre-trained checkpoint. Once we have poisoned the model, we attempt to remove the poison using CleanCLIP, which finetunes the model with a specific objective using a separate dataset. We have illustrated this in Figure 1 and summarized our key findings in Table 1.
@@ 3.2 Experimental Setup
We train a dual-encoder multimodal model on image-text paired datasets. We train models using two kinds of pre-training objectives: a) only multimodal contrastive loss ($\mathcal{L}^{pre}_{\text{MMCL}}$), and b) combination of multimodal contrastive loss and self-supervised loss in the image and text modalities ($\mathcal{L}^{pre}_{\text{MMCL}}+\mathcal{L}^{pre}_{\text{SSL}}$). Following CleanCLIP, we use a ResNet-50 as the model’s vision encoder and a transformer as the text encoder. We trained the models on two image-text paired datasets:
@@
    Conceptual Captions 3M (CC3M) (Sharma et al., 2018): This dataset has 3M image-text paired datapoints.
@@
    Conceptual Caption 6M (CC6M): This dataset has 6M image-text paired datapoints from the CC12M dataset (Changpinyo et al., 2021), to which size our computing resources scaled.
The models are trained either from scratch or finetuned from a pre-trained CLIP checkpoint (Radford et al., 2019). We train models for 64 epochs using 8 Nvidia A100 GPUs. The initial learning rate of $31e-$ 3 with cosine scheduling is used when trained from scratch and $75e-$ 7 when finetuned from a checkpoint. We use AdamW optimizer with 10,000 warmup steps (Loshchilov & Hutter, 2017). Models trained with $\mathcal{L}^{pre}_{\text{MMCL}}$ use a batch size of 256, whereas models trained with $\mathcal{L}^{pre}_{\text{MMCL}}+\mathcal{L}^{pre}_{\text{SSL}}$ use a batch size of 128. Please refer to Appendix A for the loss dynamics.
@@
Following CleanCLIP, we introduce the trigger proposed by BadNet (Gu et al., 2017) in a small subset of the training datapoints. Specifically, we add a trigger patch of size 16 $\times$ 16 sampled from a standard Gaussian at a random location in the image and subsequently change the image’s caption to be the adversary chosen label, in this case “banana”. Please see Appendix J for examples of images with trigger patch and their corresponding captions. Using the same settings as CleanCLIP, we introduce the trigger in 1,500 randomly sampled datapoints for the CC3M dataset and 3,000 randomly sampled datapoints for the CC6M dataset (a mere 0.05% of the training datapoints).
We also experiment with another kind of poisoning technique: label consistent poisoning. In this case, the trigger patch (created in the manner as mentioned above) is added to the images that have the adversary chosen label (in this case“banana”) in their captions.
Removing poison
We attempt to remove poisons from pre-trained models by finetuning them on a 100K clean image-text paired dataset using $\mathcal{L}^{pre}_{\text{MMCL}}$, $\mathcal{L}^{pre}_{\text{SSL}}$, and $\mathcal{L}^{ft}_{\text{MMCL}}+\mathcal{L}^{ft}_{\text{SSL}}$ (CleanCLIP). We consider a model to be cleaned if the ASR of that model is $\leq 5\%$.
Table 1: Key findings from our experiments: CleanCLIP is much less effective for the model where poison is induced with the stronger objective $\mathcal{L}^{pre}_{\text{MMCL}}+\mathcal{L}^{pre}_{\text{SSL}}$.
Backdoor Objective Poison Induction Change in Accuracy after Cleaning (Relative) $\uparrow$
BadNet $\mathcal{L}^{pre}_{\text{MMCL}}$ From scratch 1% gain
BadNet $\mathcal{L}^{pre}_{\text{MMCL}}$ Finetuned from Ckpt 17% loss
Label Consistent $\mathcal{L}^{pre}_{\text{MMCL}}$ From scratch 10% gain
BadNet $\mathcal{L}^{pre}_{\text{MMCL}}+\mathcal{L}^{pre}_{\text{SSL}}$ From scratch 45% loss
BadNet $\mathcal{L}^{pre}_{\text{MMCL}}+\mathcal{L}^{pre}_{\text{SSL}}$ Finetuned from Ckpt 33% loss
Label Consistent $\mathcal{L}^{pre}_{\text{MMCL}}+\mathcal{L}^{pre}_{\text{SSL}}$ From scratch 16% loss